# The large sieve for square moduli, revisited

### v3

@@ 1.1. Review of the classical large sieve
Large sieve inequalities have turned out extremely useful in analytic number theory. The classical example is the large sieve inequality for additive characters, which can be easily converted into a large sieve inequality for multiplicative characters. It originated in Linnik’s work [25] on the least quadratic non-residue modulo primes. This inequality is a very flexible tool: It can be turned into an upper bound sieve but also be used to understand the distribution of arithmetic functions in residue classes on average, for example. Its statement is as follows.
@@ 2. Additive energies of modular square roots and main result
Additive energy is an important concept in additive combinatorics. During the recent years, it has turned out to be very useful in analytic number theory too (see [8], [14], [15], [16], [22], [29], [30], for example) and will likely continue to be. This includes the work [8] by Baker, Munsch and Shparlinski on the large sieve with sparse sets of moduli. Our article builds on recent investigations of additive energies related to modular square roots. This development was started off in the paper [14] by Dunn and Zaharescu (published first on the arXiv preprint server in 2019 (arXiv:1903.03416) and appeared in 2025), where modular square roots arose in connection with bilinear forms of Salié sums. This line of investigations was continued in the paper [15] by Dunn, Kerr, Shparlinski and Zaharescu, where improvements were obtained. Our article relies on subsequent results by Kerr, Shkredov, Shparlinski and Zaharescu in [22] and [30] (see also [29] for an average result with a power saving). A concise summary of this sequence of works can be found in [14, section 1, after Theorem 1.2]. In this section, we describe the results on additive energies of modular square roots in [22] and [30], state hypotheses and formulate our main result of this article.
@@
It is conceivable that the methods used to prove the above Theorems 3 and 4 in [22] can be made work for all moduli $r$, giving results similar to (10) and (12) without assuming $r$ to be a prime. Care needs to be taken, though: Whereas for prime moduli $r$, the number of modular square roots of a given $m\bmod r$ is bounded by 2, this number may be considerably larger if $r$ is not a prime. For example, if $r=p^{2}$ is the square of a prime, then $m=0$ has precisely $p$ square roots modulo $r$, namely $\sqrt{m}=kp$ with $k=0,...,p-1$. More generally, let $r=p^{\alpha}$ be a prime power and $0\leq m=m_{1}p^{\beta}<r$ with $(m_{1},p)=1$. If $m\equiv k^{2}\bmod{r}$, then $k=k_{1}p^{\lceil\beta/2\rceil}$ for some $k_{1}\bmod{p^{\alpha-\lceil\beta/2\rceil}}$, and the congruence reduces to $m_{1}\equiv k_{1}^{2}p^{2\lceil\beta/2\rceil-\beta}\bmod{p^{\alpha-\beta}}$. If $m\not=0$, then $\beta<\alpha$ and $\beta$ is necessarily even. It follows that $m_{1}\equiv k_{1}^{2}\bmod{p^{\alpha-\beta}}$. Using Hensel’s lemma, the number of solutions $k_{1}\bmod{p^{\alpha-\beta/2}}$ of this congruence can be as large as $2p^{\beta/2}$ if $p$ is odd and $2^{\beta/2+2}$ if $p=2$. In the case when $m=0$, there are exactly $p^{\lfloor\alpha/2\rfloor}$ square roots of $m$ modulo $r$. Now using the Chinese remainder theorem, for general moduli $r$, the number of modular square roots of any $m\bmod{r}$ is bounded by $O_{\varepsilon}\left(r^{\varepsilon}(m,r)^{1/2}\right)$. It follows that the number of modular square roots modulo $r$ of all $m$ in the range $1\leq m\leq R$ is bounded by
@@ 3. Preliminaries from the theory of exponential sums
This is a consequence of [17, Lemma 2.5.]. ∎
For a convenient application of the Poisson summation formula, we will use smooth weight functions $\Phi:\mathbb{R}\rightarrow\mathbb{C}$ which fall into the Schwartz class (for details on the Schwartz class, see [32]). For a Schwartz class function as above, we define its Fourier transform $\hat{\Phi}:\mathbb{R}\rightarrow\mathbb{C}$ as
@@
$S(f,p):=\sum\limits_{\begin{subarray}{c}n\bmod{p^{m}}\\ f_{2}(n)\not\equiv 0\bmod{p}\end{subarray}}e_{p}\left(f(n)\right).$
@@
This follows from [13, inequality (1.4)]. ∎
@@
This is found in [13, Theorem 3.1]. ∎
@@
This follows from the results in [17, subsection 7.4] and [10, Theorem 1.1.5 and Lemma 1.2.1]. ∎
@@ 4.1. General approach
(17) $\sum\limits_{k=1}^{K}\left|\sum\limits_{M<n\leq M+N}a_{n}e(n\alpha_{k})\right|^{2}\ll\max\limits_{x\in\mathbb{R}}\sharp\left\{k\in\{1,...,K\}:||\alpha_{k}-x||\leq\frac{1}{N}\right\}\cdot NZ,$
where for any real number $z$, $||z||$ denotes the distance of $z$ to the nearest integer. Throughout the sequel, we will write
@@
$\Delta_{\mathcal{S}}(Q,N)=N\cdot\max\limits_{x\in[0,1]}\sharp\left\{\frac{a}{q}:q\in\mathcal{S},\ q\leq Q,\ 1\leq a\leq q,\ (a,q)=1,\ \left|\left|\frac{a}{q}-x\right|\right|\leq\Delta\right\}.$
@@ 4.2. Wolke’s treatment of prime moduli
$M:=\max\limits_{x\in[0,1]}\sharp\left\{\frac{a}{p}:p\in\mathbb{P},\ p\leq Q,\ 1\leq a\leq p-1,\ \left|\left|\frac{a}{p}-x\right|\right|\leq\Delta\right\}.$
@@
$[x-\Delta,x+\Delta]=\left[\frac{b}{r}+z-\Delta,\frac{b}{r}+z+\Delta\right]$
@@ 4.3. Application of Wolke’s approach to square moduli
Essentially, this means is that $q$ runs over a short interval of length $\delta/Q$ around $\sqrt{y}$, $m$ runs over a short interval of length $\delta rz$ around $yrz$, these two variables are linked by the congruence $m\equiv-bq^{2}\bmod{r}$, and the count of pairs $(m,q)$ in the resulting small box is averaged over the variable $y$ of size $Q^{2}$. This average is scaled by a factor of $\delta^{-1}$. Note that with increasing $\delta$, the area of the box increases by a factor of $\delta^{2}$. So it appears that taking $\delta$ as small as possible in (24) is optimal. However, after Fourier analytic transformations, it may be better to choose $\delta$ differently. So it helps to have flexibility in the choice of $\delta$. Writing the congruence in the form $q^{2}\equiv-\overline{b}m\bmod{r}$, the task becomes to count squares in small segments of residue classes modulo $r$, which is a situation similar to that considered by Wolke for primes.
@@ 5.1. Direct count
Following the method in [1], it turns out that the averaging over $y$ allows one to remove the term "$1+$" in "$1+\delta rz$" above. Consequently, we get
@@ 5.3. Advantages and disadvantages
A direct count is simple but unable to avoid the contribution "$1+$" in the term
@@ 7.1. Sketch for our particular situation
(46) $|\Sigma|^{2}\ll Q\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ \left|\left|(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r\right|\right|\ll Q^{-1/2}\end{subarray}}1.$
The same can be achieved via an application of the double large sieve, Proposition 12, with the convention that the modular square root $\sqrt{jm}$ lies in the range $0\leq\sqrt{jm}<r$ and so $10\leq\sqrt{jm}/r<$ 1. Discarding the summation condition $\left|\left|(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r\right|\right|\ll Q^{-1/2}$, we get the trivial bound $|\Sigma|^{2}\ll Q^{2}$. Hence, we need to make use of this condition to beat the trivial bound. Heuristically, we expect that the "event" $\left|\left|(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r\right|\right|\ll Q^{-1/2}$ occurs with a probability of $Q^{-1/2}$ when picking $m_{1}$ and $m_{2}$ randomly. Trivially, this event occurs when $m_{1}=m_{2}$. So it is reasonable to conjecture that
@@
$\Sigma^{\prime}=\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ \left|\left|(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r\right|\right|\ll Q^{-1/2}\end{subarray}}1=\sum\limits_{\begin{subarray}{c}\lambda\in\mathbb{Z}\\ |\lambda|\ll Q^{-1/2}r\end{subarray}}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ \sqrt{jm_{1}}-\sqrt{jm_{2}}\equiv\lambda\bmod{r}\end{subarray}}1$
@@
in our situation. (Recall that the trivial bound is $P(x)\ll Q^{1/2}$.) If the equations in (39) hold and $r$ is smaller than $Q^{3/2}$, then the analytic term (43) oscillates, so removing it by partial summation comes with a cost. Proceeding along the same lines as above, it turns out that these costs are moderate enough to still get a non-trivial bound under the Hypothesis 6 if $Q=N^{1/3}$ and $Q^{1+\eta}\leq r\leq Q^{3/2}$, where $\eta>0$ is arbitrarily small but fixed. In the next subsection, we will work out the details of these calculations for large $r$’s and general $Q$, $N$ and $z$.
Remark. If one avoids Wolke’s approach using Diophantine approximation by fractions $b/r$, then one needs to bound the number of Farey fractions $a/q^{2}$ with $q\sim Q$ in intervals of the form $[a_{0}/q_{0}^{2}-\Delta,a_{0}/q_{0}^{2}+\Delta]$, where $q_{0}\sim Q$ is fixed. One may approach this problem in the same way as above via modular square roots modulo $q_{0}^{2}$. However, calculations show that in this case, our hypotheses on additive energies just lead us to a bound of $\ll Q^{1/2}$ for the said number of Farey fractions if $Q^{3}=N$, and therefore, no progress has been made along these lines. This indicates that an initial Diophantine approximation by Farey fractions $b/r$ is indeed essential. It allows us to replace the fractions $a_{0}/q_{0}^{2}$ by fractions $b/r$ with much smaller denominators.
@@ 7.2. Conditional treatment of the general case
where we take the rapid decay of $W$ into account. Further, since $\Phi_{5}$ has compact support $[C_{0},C_{1}]$ for suitable constants $C_{1}>C_{0}>0$, the $m$-range can be restricted to $M_{0}\leq m\leq M_{1}$ with
@@
By convention, we take the modular square root of $jm$ always in such a way that $10\leq\sqrt{jm}/r<$ 1. To bound the double sum above, we now apply the double large sieve, Proposition 12. Here we let the $\alpha_{k}$’s run over all integers with modulus not exceeding $L_{1}$ and the $\beta_{l}$’s over all real numbers $(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r$ arising from pairs of integers $(m_{1},m_{2})$ with $M_{0}\leq m_{1},m_{2}\leq M$. Taking $A:=L_{1}$ and $B:=1$ and setting all coefficients equal to 1, Proposition 12 yields
@@ 8.1. Sketch for our particular situation
$\begin{split}\left(\sum\limits_{|l|\ll Q^{1/2}/H}|\Sigma^{\prime}(h,l)|\right)^{2}=&\left(\sum\limits_{|l|\ll Q^{1/2}/H}\epsilon_{h,l}\Sigma^{\prime}(h,l)\right)^{2}\\ \ll&\frac{Q}{H^{2}}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ ||(f_{j,h}(m_{1})-f_{j,h}(m_{2}))/r||\ll H/Q^{1/2}\end{subarray}}1,\end{split}$
@@
$\begin{split}\bigg{(}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ ||(f_{j,h}(m_{1})-f_{j,h}(m_{2}))/r||\leq H/Q^{1/2}\end{subarray}}1\bigg{)}^{2}=&\bigg{(}\sum\limits_{|\lambda|\ll rH/Q^{1/2}}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ f_{j,h}(m_{1})-f_{j,h}(m_{2})\equiv\lambda\bmod{r}\end{subarray}}1\bigg{)}^{2}\\ \ll&\frac{rH}{Q^{1/2}}\cdot\sum\limits_{\lambda\bmod r}\bigg{(}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ f_{j,h}(m_{1})-f_{j,h}(m_{2})\equiv\lambda\bmod{r}\end{subarray}}1\bigg{)}^{2}\\ \ll&\frac{rH}{Q^{1/2}}\cdot F_{2}(cQ^{1/2};j,h,r)\\ =&HQ^{1/2}F_{2}(cQ^{1/2};j,h,r),\end{split}$
@@
$\begin{split}\sum\limits_{\begin{subarray}{c}m_{1},m_{2}\asymp Q^{1/2}\\ ||(f_{j,h}(m_{1})-f_{j,h}(m_{2}))/r||\leq H/Q^{1/2}\end{subarray}}1\ll H^{1/2}Q^{1/4}\left((h,r)\cdot\frac{Q^{2}}{r}+Q\right)^{1/2}N^{\varepsilon}\ll H^{1/2}Q^{3/4}N^{\varepsilon}(h,r)^{1/2},\end{split}$
@@
In the next subsection, we will work out the details of the above calculations for medium-sized $r$’s and general $Q$, $N$ and $z$. It will turn out that if the equations in (39) hold, we obtain a non-trivial bound for $P(x)$ if $Q^{1/2+\eta}\leq r\leq Q^{9/8}$ along these lines.
@@ 8.2. Conditional treatment of the general case
We note that like $\Phi_{5}$, this function $V_{h}$ has compact support in $[C_{0},C_{1}]\subset\mathbb{R}_{>0}$. We remove the weight
@@ 9.1. Sketch for our particular situation
Since $V_{h}$ has compact support in $[C_{0},C_{1}]\subset\mathbb{R}_{>0}$ and $hQ^{-1/2}\ll 1$, we have
@@
This is fine if $\varepsilon$ is small enough, and once again, we have obtained a non-trivially bound for $P(x)$ this time at the point $r=Q^{1/2}$. Again, we will carry out the above method for general $Q$, $N$ and $z$ in the following subsection. In the situation of (39), we will obtain a non-trivial bound for $P(x)$ if $Q^{\eta}\leq r\leq Q^{15/26-\eta}$ for some $\eta>0$.

### v4

@@ 1.1. Review of the classical large sieve
Large sieve inequalities have turned out extremely useful in analytic number theory. The classical example is the large sieve inequality for additive characters, which can be easily converted into a large sieve inequality for multiplicative characters. It originated in Linnik’s work [26] on the least quadratic non-residue modulo primes. This inequality is a very flexible tool: It can be turned into an upper bound sieve but also be used to understand the distribution of arithmetic functions in residue classes on average, for example. Its statement is as follows (see [11, Satz 5.2.2] for the specific case of Farey fractions).
@@ 2. Additive energies of modular square roots and main result
Additive energy is an important concept in additive combinatorics. During the recent years, it has turned out to be very useful in analytic number theory too (see [8], [15], [16], [17], [23], [30], [31], for example) and will likely continue to be. This includes the work [8] by Baker, Munsch and Shparlinski on the large sieve with sparse sets of moduli. Our article builds on recent investigations of additive energies related to modular square roots. This development was started off in the paper [15] by Dunn and Zaharescu (published first on the arXiv preprint server in 2019 (arXiv:1903.03416) and appeared in 2025), where modular square roots arose in connection with bilinear forms of Salié sums. This line of investigations was continued in the paper [16] by Dunn, Kerr, Shparlinski and Zaharescu, where improvements were obtained. Our article relies on subsequent results by Kerr, Shkredov, Shparlinski and Zaharescu in [23] and [31] (see also [30] for an average result with a power saving). A concise summary of this sequence of works can be found in [15, section 1, after Theorem 1.2]. In this section, we describe the results on additive energies of modular square roots in [23] and [31], state hypotheses and formulate our main result of this article.
@@
It is conceivable that the methods used to prove the above Theorems 3 and 4 in [23] can be made work for all moduli $r$, giving results similar to (10) and (12) without assuming $r$ to be a prime. Care needs to be taken, though: Whereas for prime moduli $r$, the number of modular square roots of a given $m\bmod r$ is bounded by 2, this number may be considerably larger if $r$ is not a prime. For example, if $r=p^{2}$ is the square of a prime, then $m=0$ has precisely $p$ square roots modulo $r$, namely $\sqrt{m}=kp$ with k=0,,p1k=0,...,p-1. More generally, let $r=p^{\alpha}$ be a prime power and $0\leq m=m_{1}p^{\beta}<r$ with $(m_{1},p)=1$. If $m\equiv k^{2}\bmod{r}$, then $k=k_{1}p^{\lceil\beta/2\rceil}$ for some $k_{1}\bmod{p^{\alpha-\lceil\beta/2\rceil}}$, and the congruence reduces to $m_{1}\equiv k_{1}^{2}p^{2\lceil\beta/2\rceil-\beta}\bmod{p^{\alpha-\beta}}$. If $m\not=0$, then $\beta<\alpha$ and $\beta$ is necessarily even. It follows that $m_{1}\equiv k_{1}^{2}\bmod{p^{\alpha-\beta}}$. Using Hensel’s lemma, the number of solutions $k_{1}\bmod{p^{\alpha-\beta/2}}$ of this congruence can be as large as $2p^{\beta/2}$ if $p$ is odd and $2^{\beta/2+2}$ if $p=2$. In the case when $m=0$, there are exactly $p^{\lfloor\alpha/2\rfloor}$ square roots of $m$ modulo $r$. Now using the Chinese remainder theorem, for general moduli $r$, the number of modular square roots of any $m\bmod{r}$ is bounded by $O_{\varepsilon}\left(r^{\varepsilon}(m,r)^{1/2}\right)$. It follows that the number of modular square roots modulo $r$ of all $m$ in the range $1\leq m\leq R$ is bounded by
@@
Acknowledgements. The author wishes to thank the anonymous referee for carefully going through this article and the Ramakrishna Mission Vivekananda Educational and Research Institute for providing excellent working conditions.
@@ 3. Preliminaries from the theory of exponential sums
This is a consequence of [18, Lemma 2.5.]. ∎
For a convenient application of the Poisson summation formula, we will use smooth weight functions $\Phi:\mathbb{R}\rightarrow\mathbb{C}$ which fall into the Schwartz class (for details on the Schwartz class, see [33]). For a Schwartz class function as above, we define its Fourier transform $\hat{\Phi}$$\mathbb{R}\rightarrow\mathbb{C}$ as
@@
$S(f,p):=\sum\limits_{\begin{subarray}{c}n\bmod{p}\\ f_{2}(n)\not\equiv 0\bmod{p}\end{subarray}}e_{p}\left(f(n)\right).$
@@
This follows from [14, inequality (1.4)]. ∎
@@
This is found in [14, Theorem 3.1]. ∎
@@
This follows from the results in [18, subsection 7.4] and [10, Theorem 1.1.5 and Lemma 1.2.1]. ∎
@@ 4.1. General approach
[…]
where for any real number $z$, $z||$ denotes the distance of $z$ to the nearest integer. Throughout the sequel, we will write
@@
[…]
@@ 4.2. Wolke’s treatment of prime moduli
[…]
@@
$x-\Delta,x+\Delta]=\left[\frac{b}{r}+z-\Delta,\frac{b}{r}+z+\Delta\right]$
@@ 4.3. Application of Wolke’s approach to square moduli
Essentially, this means is that $q$ runs over a short interval of length $\delta/Q$ around $\sqrt{y}$, $m$ runs over a short interval of length $\delta rz$ around yrzyrz, these two variables are linked by the congruence $m\equiv-bq^{2}\bmod{r}$, and the count of pairs $(m,q)$ in the resulting small box is averaged over the variable $y$ of size $Q^{2}$. This average is scaled by a factor of $\delta^{-1}$. Note that with increasing $\delta$, the area of the box increases by a factor of $\delta^{2}$. So it appears that taking $\delta$ as small as possible in (24) is optimal. However, after Fourier analytic transformations, it may be better to choose $\delta$ differently. So it helps to have flexibility in the choice of $\delta$. Writing the congruence in the form $q^{2}\equiv-\overline{b}m\bmod{r}$, the task becomes to count squares in small segments of residue classes modulo $r$, which is a situation similar to that considered by Wolke for primes.
@@ 5.1. Direct count
Following the method in [1], it turns out that the averaging over $y$ allows one to remove the term "1+1+" in $1+\delta rz$" above. Consequently, we get
@@ 5.3. Advantages and disadvantages
A direct count is simple but unable to avoid the contribution "1+1+" in the term
@@ 7.1. Sketch for our particular situation
[…]
@@
[…]
@@
in our situation. (Recall that the trivial bound is $P(x)\ll Q^{1/2}$.) If the equations in (39) hold and $r$ is smaller than $Q^{3/2}$, then the analytic term (43) oscillates, so removing it by partial summation comes with a cost. Proceeding along the same lines as above, it turns out that these costs are moderate enough to still get a non-trivial bound under the Hypothesis 6 if $Q=N^{1/3}$ and $Q^{1+\eta}\leq r\leq Q^{3/2}$, where $\eta>0$ is arbitrarily small but fixed. In the next subsection, we will work out the details of these calculations for large $r$’s and general $Q$, NN and $z$.
Remark. If one avoids Wolke’s approach using Diophantine approximation by fractions $b/r$, then one needs to bound the number of Farey fractions $a/q^{2}$ with $q\sim Q$ in intervals of the form $a_{0}/q_{0}^{2}-\Delta,a_{0}/q_{0}^{2}+\Delta]$, where $q_{0}\sim Q$ is fixed. One may approach this problem in the same way as above via modular square roots modulo $q_{0}^{2}$. However, calculations show that in this case, our hypotheses on additive energies just lead us to a bound of $\ll Q^{1/2}$ for the said number of Farey fractions if $Q^{3}=N$, and therefore, no progress has been made along these lines. This indicates that an initial Diophantine approximation by Farey fractions $b/r$ is indeed essential. It allows us to replace the fractions $a_{0}/q_{0}^{2}$ by fractions $b/r$ with much smaller denominators.
@@ 7.2. Conditional treatment of the general case
where we take the rapid decay of $W$ into account. Further, since $\Phi_{5}$ has compact support $C_{0},C_{1}]$ for suitable constants $C_{1}>C_{0}>0$, the $m$-range can be restricted to $M_{0}\leq m\leq M_{1}$ with
@@
By convention, we take the modular square root of $jm$ always in such a way that $0\leq\sqrt{jm}/r<1$. To bound the double sum above, we now apply the double large sieve, Proposition 12. Here we let the $\alpha_{k}$’s run over all integers with modulus not exceeding $L_{1}$ and the $\beta_{l}$’s over all real numbers $(\sqrt{jm_{1}}-\sqrt{jm_{2}})/r$ arising from pairs of integers $(m_{1},m_{2})$ with $M_{0}\leq m_{1},m_{2}\leq M$. Taking $A:=L_{1}$ and $B:=1$ and setting all coefficients equal to 1, Proposition 12 yields
@@ 8.1. Sketch for our particular situation
[…]
@@
[…]
@@
[…]
@@
In the next subsection, we will work out the details of the above calculations for medium-sized $r$’s and general $Q$, NN and $z$. It will turn out that if the equations in (39) hold, we obtain a non-trivial bound for $P(x)$ if $Q^{1/2+\eta}\leq r\leq Q^{9/8}$ along these lines.
@@ 8.2. Conditional treatment of the general case
We note that like $\Phi_{5}$, this function $V_{h}$ has compact support in $C_{0},C_{1}]\subset\mathbb{R}_{>0}$. We remove the weight
@@ 9.1. Sketch for our particular situation
Since $V_{h}$ has compact support in $C_{0},C_{1}]\subset\mathbb{R}_{>0}$ and $hQ^{-1/2}\ll 1$, we have
@@
This is fine if $\varepsilon$ is small enough, and once again, we have obtained a non-trivially bound for $P(x)$ - this time at the point $r=Q^{1/2}$. Again, we will carry out the above method for general $Q$, NN and $z$ in the following subsection. In the situation of (39), we will obtain a non-trivial bound for $P(x)$ if $Q^{\eta}\leq r\leq Q^{15/26-\eta}$ for some $\eta>0$.